OpenStudy (anonymous):

Find the Ph of a solution whose pOH is 1.36

OpenStudy (anonymous):

Kw = [H^+][OH-] = 1.00 x 10^-14, pH + pOH = 14, pH = 14.00 - pOH pH = 14.00 - 1.36 = 12.64